OpenStudy (anonymous):

find the intercepts of f(x)=1/2+2 and g(x)=2-1/2x

OpenStudy (anonymous):

is f(x) written correctly? and is g(x) = 2 - 1/(2x) or g(x) = 2 - (1/2)x ??